OpenStudy (anonymous):

What property is shown by the equation? Column A Column B 1. –24 • 0 = 0 2. 3. –18 • (–1) = 18 4. 5 ÷ (–1) = –5 A. Reciprocal Property of Multiplication B. Identity Property of Division C. Negative One Property of Multiplication D. Zero Property of Multiplication E. Negative One Property of Division

OpenStudy (anonymous):

@chucky00

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

-24 * 0 = 0 Do you know what property that is? It's pretty easy to guess.

OpenStudy (igreen):

In -24 * 0 you're multiplying 0. In 18 * (-1) =

OpenStudy (anonymous):

d and e

OpenStudy (anonymous):

http://openstudy.com/updates/54526c55e4b0be646b9a24af i will givemetalz

OpenStudy (igreen):

In 18 * (-1) = -18 you're multiplying -1. (KEY WORDS: -1 and MULTIPLY) In 5 / (-1) = -5 you're dividing -1. (KEY WORDS: -1 and DIVIDE) THESE ARE EASY! The properties tell you..

OpenStudy (igreen):

24 * 0 = 0 You're multiplying 0 (KEY WORDS: 0 and MULTIPLY) This is the ZERO property of MULTIPLICATION. In 18 * (-1) = -18 you're multiplying -1. (KEY WORDS: -1 and MULTIPLY) This is the NEGATIVE ONE property of MULTIPLICATION. In 5 / (-1) = -5 you're dividing -1. (KEY WORDS: -1 and DIVIDE) This is the NEGATIVE ONE property of DIVISION. Do you understand?

OpenStudy (anonymous):

thanks